17. Which of the ordered pairs
(–3, 1), (1, 0), (–11, –3), (–27, 7)
are solutions for the equation x + 4y = 1?

Answer by josmiceli(19441)

I know it can't be (-11, -3). That would mean the sum of two negative numbers is a positive- not true.
Test the others
(-3) + 4(1) = 1
TRUE
(1) + 4(0) = 1
TRUE
(-27) + 4(7) = 1
-27 + 28 = 1
TRUE